OpenStudy (anonymous):

simplify (sqrt21+9sqrt10)(sqrt6-4sqrt35)

OpenStudy (anonymous):

\[(\sqrt{21} + 9\sqrt{10})(\sqrt{6} - 4\sqrt{35})\]

mathslover (mathslover):

can u expand it ?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

Do you know what the distributive property is?

OpenStudy (anonymous):

yes

mathslover (mathslover):

\[\large{(\sqrt{21}+9\sqrt{10})(\sqrt{6}-4\sqrt{35})}\] \[\large{\sqrt{21*6}-4\sqrt{21*35}+9\sqrt{10*6}-36\sqrt{10*35}}\]

OpenStudy (anonymous):

thxs i remember this now